# pnasresearcharticle

### v1

@@ Abstract
The preexisting MSEβ inversion within the lower free troposphere is typically marked by a pronounced temperature inversion or a layer where the temperature decreases most gradually with height. While this study proposed the theoretical framework based on evolution of local convective instability, the formation of preexisting inversion layers might be non-local due to advection of a warm layer from the upstream region or the adiabatically warmed air from the subsidence of large-scale circulation (e.g., anticyclones). Thus, it might be necessary to investigate the origins and trajectories of the inversion layers from the Lagrangian perspective (10, 11, 16) to provide insight into the formation or potential magnitude of MSE, and thus improve the predictability of local potential moist heat and convection. This is also important in understanding the role of preexisting inversion layers in future climate, as the shift in land types or non-uniform warming by regions may affect the formation and transportation trajectories of inversion layers from region to region (17, 11).
[…]
Buoyancy and temperature perturbation
@@
Considering an undiluted air parcel adiabatically lifted upward from the near-surface (2 m above ground surface; denoted by subscript ββ in derivations) and neglecting the relatively small contribution to density changes due to pressure perturbation and virtual temperature correction(18, 19), the parcel buoyancy at a given height is proportional to differences in sensible heat between the air parcel and the environment (), where is temperature perturbations of the air parcel (, with subscript ββ referring to the air parcel) with respect to the environment () and J kg-1 K-1 is the the specific heat capacity of air at constant pressure. During the adiabatic process, it is common to assume that the air parcel conserves its moist static energy (MSE) such that . is defined as the sum of sensible heat (), latent heat (), and geopotential energy (), where J kg-1 is the latent heat of vaporization, m s-2 is gravitational acceleration, is temperature, is specific humidity, and is height above sea level. Therefore, the vertical profile of is given by
where is dry static energy of the environment.
To derive solely as a function of environmental parameters, independent of the lifted air parcelβs profiles, the specific humidity of the air parcel () in Eq.5 can be further approximated using environmental parameters that vary with height () relative to the lifted condensation level (LCL).
@@
For , the air parcel is saturated, such that the parcel specific humidity equals its saturated specific humidity (, with superscript ββ referring to a quantity at saturation), which is a function of the parcel temperature and air pressure. As pressure difference between the air parcel and environment is negligible (18, 19), the difference in saturated specific humidity between the air parcel and environment is caused by difference in temperature, and thus can be written as a linear relation to the environmental saturated specific humidity () via . We further assume constant and approximate it by using Clausius-Clapeyron equation, where J kg-1 K-1 is gas constant for water vapor. Hence,
Substituting Eqs.6 and 7 into Eq.5 and rearranging it gives
Eq.8 well recovers the temperature perturbation profiles ( , Fig. S3). The denominator is always positive, ranging from 3 in the lower free troposphere above LCL to 1 in the upper free troposphere ( , Fig. S3). To simplify and derive a more intuitive form, we approximate the denominator as 1, which does not alter the sign, vertical structure, or relative magnitude of above LCL but only slightly amplifies temperature difference near LCL ( , Fig. S3). Therefore,
Eq.9 provides an alternative approach to quantify temperature perturbation profile from the MSE prospective (a more precise form is given by Eq.8; , Fig. S3). This eliminates the need to first calculate the temperature profile of a hypothetically lifted air parcel. Instead, the vertical profile of parcel buoyancy can be determined directly by comparing the parcelβs initial MSE (i.e., ) with the environmental static energy profile ( if or if ). Furthermore, this approach is equivalent to comparing near-surface potential temperature with the potential temperature profile for or the near-surface equivalent potential temperature with the saturation equivalent potential temperature profile for .
@@
Next, we derive a scaling for CAPE from Eq.10 combining simple assumptions. Since is zero at both LFC and EL and reaches its maximum approximately midway between these levels ( 500 hPa, typically above LCL), CAPE can be geometrically approximated as the area of a triangle with being the base and at 500 hPa (, with superscript ββ referring to a quantity at 500 hPa) being the height. Further neglecting virtual temperature correction (i.e., ) and substituting Eq.8, we have that
Here we focus on midlatitude extreme moist heat conditions that CAPE is in general high (1000 J kg-1). For those cases, the LFC is near the lower free troposphere and EL is close to the tropopause, and we approximate them by hPa and hPa. The mean K gives kg kg-1. Substituting these typical values into Eq.11, we provide a scaling for midlatitude high CAPE:
Previous studies have explored the relationship between CAPE and environmental MSE, either under the radiativeβconvective equilibrium framework (22, 23, 12) or through linear regression methods (24). In contrast, our derivation begins with the fundamental definitions of parcel buoyancy and CAPE, similar to (25) but incorporating minimal assumptions about the geometric characteristics of CAPE, to establish a robust linear relationship between CAPE and key atmospheric parameters near the surface and in the free troposphere. The coefficient in Eq.12 may vary with , , and , which can be precisely adjusted using the full scaling equation (Eq.11) when applied to different regions or datasets. Here, the derived coefficient of 0.22 for high-CAPE environments over midlatitude land aligns closely with the linear regression results (0.22β0.29) reported for high CAPE over continental North America (24).
This research is part of the MIT Climate Grand Challenge on Weather and Climate Extremes. Support was provided by Schmidt Sciences, LLC. We thank Kerry Emanuel, Paul OβGorman, P. J. Tuckman, and Divya Rea for helpful discussions.
[…]

### v3

@@ Abstract
The preexisting MSE∗ inversion within the lower free troposphere is typically marked by a pronounced temperature inversion or a layer where the temperature decreases most gradually with height (61, 62, 63, 15, 64, 65, 55, 56, 66). While this study proposed the theoretical framework based on evolution of local convective instability, the formation of preexisting inversion layers might be non-local due to advection of a warm layer from an adjacent region or the adiabatically warmed air from the subsidence of large-scale circulation (e.g., anticyclones) (92, 56, 61, 65, 15). Thus, it might be worthwhile to investigate the origins and evolutions of the inversion layers from the Lagrangian perspective (76, 93, 94, 55, 74, 56) to provide insight into the formation and potential magnitude of MSE, and thus improve the predictability of local potential moist heat and convection. This is also critical for understanding the role of preexisting lower free tropospheric inversion layers in future climates, as shifts in land cover or regionally uneven warming could affect the formation and transport pathways of inversion layers between regions (95, 96, 97, 98, 66, 53, 56), thereby potentially shaping future spatial patterns of extreme heat and convection (99, 100, 101, 26).
Data
We use 3-hourly near-surface and pressure-level ERA5 reanalysis data with a horizontal grid spacing of 0.5∘ latitude by 0.5∘ longitude during 1980–2022 (36, 37), for land grid points between 35∘N and 75∘N at elevations lower than 1000 m (20854 grid points per year in total). ERA5 reanalysis data are widely used in studies of global temperature extremes and convective environments, as they accurately capture observed temperature and moisture profiles in general (102, 103, 56), including inversion features within the lower free troposphere (64, 66). In this study, calculations for all variables are initially performed at the 3-hourly interval and then sampled based on the daily maximum MSEs to create a dataset of daily maximum moist heat. Our analyses focus on the annual maximum moist heat, obtained by further sampling the dataset given the annual maximum MSEs at each grid point for each year. This process yields 43 samples per land grid point, corresponding to the maximum MSEs for each year during 1980–2022. For analyses related to extreme near-surface dry heat, the data are generated in the same manner but are conditioned on the maximum near-surface dry heat, defined as the maximum 2-m air temperature (Ts).
MSE and WBT
MSE is defined as the sum of sensible heat (), latent heat (), and geopotential energy (), i.e., , where J kg-1 K-1 is the the specific heat capacity of air at constant pressure, J kg-1 is the latent heat of vaporization, m s-2 is gravitational acceleration, is air temperature, is specific humidity, and is height above sea level (47). Near-surface (denoted by subscript “”) MSE, , is calculated using quantities at 2-m above ground surface, where the geopotential height equals to 2 plus the surface elevation in meters. Saturated (denoted by superscript “”) MSE, , is calculated by replacing with the saturated specific humidity (), which is a function of air temperature and pressure. In an adiabatic process, MSE is generally considered a conserved quantity given hydrostatic balance (47, 104, 105), interconvertible with equivalent potential temperature (106, 107, 35). As a part of MSE, dry static energy (DSE) is the sum of and , which is conserved under dry adiabatic conditions and is thermodynamically equivalent to potential temperature (47, 108). is identified as the maximum MSE∗ within the lower free troposphere, specifically at or above but below 300 hPa.
is calculated based on the conversation of (109, 16). Specifically, for each , we solve the equation for the corresponding , where is the saturated specific humidity at temperature and surface pressure . As is proportional with based on the Clausius–Clapeyron scaling (110), exhibits a nearly one-to-one relationship with (Fig. S1) (16, 15). There is a small variation in WBTs given a value of MSEs, due to dependence on the elevation variance (0–1000 m; Fig. S1). Similarly, we calculate the maximum potential (Fig. 4) by solving from the equation at the time of annual maximum .
Composite time series
Composite time series analyses (Fig. 3–) are based on a subset of extreme moist heat and convection cases. To focus on the locally most extreme moist cases, we first select cases of annual maximum when its CAPEc corresponds to the local annual maximum CAPE (i.e., cases with moist heat and convection both maximized at the same time during the year). This step on average yields 9270 cases per year (i.e., 44.5% of all cases), with mean CAPE of up to 3011 J kg-1 and mean CIN as low as 12 J kg-1. There are a few cases over far northern Canada (Nunavut area), where convection is rare and the annual maximum CAPE is less than 500 J kg-1. Thus, those cases with CAPE 500 J kg-1 are excluded, which further yields 8959 cases per year. Composite time series based on this large group of cases is shown in Fig. S4. The composite time series indicates a substantial increase of MSEs bounded by MSE (Fig. S4) and the buildup of CAPE bounded by the potential CAPE (Fig. S4), along with the removal of the most negative buoyancy (Fig. S4). These results are broadly consistent with that shown in main text (Fig. 3–) .
We further select cases where the theoretical maximum potential intensities (Eqs.3–4) are likely achieved and the preexisting energy barrier is relatively large, by conditioning on CIN 5 J kg-1 and the maximum CIN of at least 50 J kg-1 within 7 days preceding the critical time. This leads to 2502 cases per year (i.e., 12% of all cases), with the composite time series shown in the main text (Fig. 3–).
Parcel buoyancy and temperature perturbation
@@
Considering an undiluted air parcel adiabatically lifted upward from the near-surface (2 m above ground surface; denoted by subscript “” in derivations) and neglecting the relatively small contribution to density changes due to pressure perturbation and virtual temperature correction(111, 47), the parcel buoyancy at a given height () is proportional to differences in sensible heat between the air parcel and the environment (), where is temperature perturbations of the air parcel (, with subscript “” referring to the air parcel) with respect to the environment () and J kg-1 K-1 is the the specific heat capacity of air at constant pressure. During an adiabatic process, it is common to assume that the air parcel conserves its moist static energy such that . Therefore, the vertical profile of or is given by
where is vertical profile of the environmental dry static energy.
To derive parcel buoyancy solely as a function of environmental parameters, independent of the lifted air parcel’s profiles, the specific humidity of the air parcel () in Eq.5 can be further approximated using environmental parameters that vary with height () relative to the lifted condensation level (LCL).
@@
For , the air parcel is saturated, such that the parcel specific humidity equals its saturated specific humidity (, with superscript “” referring to a quantity at saturation), which is a function of the parcel temperature and air pressure. As pressure difference between the air parcel and environment is negligible (111, 47), the difference in saturated specific humidity between the air parcel and environment is caused by difference in temperature, and thus can be written as a linear relation to the environmental saturated specific humidity () via . We further assume constant and approximate it by using Clausius-Clapeyron equation, where J kg-1 K-1 is gas constant for water vapor. Hence,
Eq.8 recovers the vertical profiles of temperature perturbations (Fig. S6). The denominator is always positive, ranging from 3 in the lower free troposphere above the LCL to 1 in the upper free troposphere (Fig. S6), which does not alter the sign of buoyancy across the atmospheric column or the location of the most negative buoyancy within the lower free troposphere (Fig. S6). Therefore, Eq.8 is further simplified in a scaling format,
Eq.9 provides an alternative approach to quantify temperature perturbation profile from the MSE prospective (a more precise form is given by Eq.8; Fig. S6). This eliminates the need to first calculate the temperature profile of a hypothetically lifted air parcel. Instead, the vertical profile of parcel buoyancy can be determined directly by comparing the parcel’s initial MSE (i.e., ) with the environmental static energy profile ( if or if ). Furthermore, this approach is equivalent to comparing near-surface potential temperature with the potential temperature profile for or the near-surface equivalent potential temperature with the saturation equivalent potential temperature profile for .
@@
Next, we derive a scaling for CAPE from Eq.10 combining simple assumptions. Since is zero at both LFC and EL and reaches its maximum approximately midway between these levels ( 500 hPa, typically above LCL at the height of minimum MSE∗; Fig. S3), CAPE can be geometrically approximated as the area of a triangle with being the base and at 500 hPa (, with superscript “” referring to a quantity at 500 hPa) being the height. Further neglecting virtual temperature correction (i.e., ) and substituting Eq.8, we have that
Here we focus on midlatitude extreme moist heat conditions where CAPE is in general high (1000 J kg-1). For those cases, the LFC is near the lower free troposphere and EL is close to the tropopause, and we approximate them by hPa and hPa. The mean K gives kg kg-1. Substituting these typical values into Eq.11, we provide a scaling for midlatitude high CAPE:
Previous studies have explored the relationship between CAPE and environmental MSE, either under the radiative–convective equilibrium framework (114, 115, 52) or through linear regression methods (116). In contrast, our derivation begins with the fundamental definitions of parcel buoyancy and CAPE, similar to (117) but incorporating minimal assumptions about the geometric characteristics of CAPE, to establish a robust linear relationship between CAPE and key atmospheric parameters near the surface and in the free troposphere. The coefficient in Eq.12 may vary with , , and , which can be precisely adjusted using the full scaling equation (Eq.11) when applied to different regions or datasets. Here, the derived coefficient of 0.22 for high-CAPE environments over midlatitude land aligns closely with the linear regression results (0.22–0.29) reported for high CAPE over continental North America (116).
Data Availability
The 3-hourly ERA5 reanalysis data (36, 37) from 1980–2022 are publicly available at https://rda.ucar.edu/datasets/ds633.0/.
Code Availability
The xcape python package is available at https://github.com/xgcm/xcape/tree/master. The Metpy python package is available at https://unidata.github.io/MetPy/latest/index.html.
This research is part of the MIT Climate Grand Challenge on Weather and Climate Extremes. Support was provided by Schmidt Sciences, LLC. We thank Kerry Emanuel, Paul O’Gorman, P. J. Tuckman, and Divya Rea for helpful discussions.